ABC Pvt Ltd proposes to buy a truck that costs 𝑅𝑀 50,000. The company has two alternatives: Alternative I: Buy from Nice Car Ltd by making a down payment of 𝑅𝑀 15,000 and settling the balance with 60 monthly payments at 10 % per annum flat rate. Alternative II: Buy from Cheap Car Ltd by making a down payment of 𝑅𝑀 10,000 and settling the balance with 60 monthly payments of 𝑅𝑀 690 each. Which alternative should the company select? Show it by calculation in detail for both alternatives.Β 


1
Expert's answer
2021-02-24T07:39:55-0500

The company should choose alternative 2 because it has low interest of Rm 1200.

The first alternative interest

"=5000-15000=35000"

"=60*0.01\/12*35000=17500"

"=35000+17500=52500"

Interest will be

"=52500-5000"

"=Rm 2500"

The second alternative interest

"=60*690"

"=Rm41400"

Interest will be

"=50000-10000=40000"

"=41400-40000= Rm 1400"